OpenStudy (anonymous):

solve by the elimination method 5u+3v=-17, 2u+v=-7 would there be a solution or not?

OpenStudy (anonymous):

There is definitely a solution to this system of equations. Solving this by elimination, you can multiply the second equation by -3: 5u + 3v = -17 -6u - 3v = 21 Now add the second equation to the first: (5 - 6)u + (3 - 3)v = 21 - 17 -u = 4 -> u = -4 Now, just put that value for "u" into either of the original equations. Using the first equation: 5(-4) + 3v = -17 -20 + 3v = -17 3v = 3 -> v = 1 And now you have your values for both "u" and "v" and you're done!

OpenStudy (anonymous):

All good now, @tootlescar47 ?

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @tootlescar47